# A Multi-agent LLM-based JUnit Test Generation with Strong Oracles

### v2

@@ 6.2 Threats to validity
Construct Validity. We evaluate CANDOR using code coverage, mutation score, and oracle correctness. While real bug detection on datasets like Defects4J is another useful metric, we did not include it because such project-level datasets are beyond the scope of this work, as discussed earlier. Instead, we use mutation score as a proxy for bug-finding capability.
@@
Internal Validity. Our results may be influenced by the choice of LLMs used in the framework. We use LLaMA 3.1 70B as the basic LLM and DeepSeek R1 as the reasoning LLM, selected based on preliminary experiments for their stability and open-source availability. While other popular models like GPT, Grok, and Mistral exist, exhaustively evaluating all LLMs is beyond the scope of this work. The choice of LLMs may affect performance, but we aim to demonstrate the potential effectiveness of our method rather than benchmark specific models, which might indeed yield even better results. Another potential threat is data leakage, wherein our datasets may exist in the pretraining data of the LLMs, leading to inflated performance. To mitigate this, we evaluate mutation score, which relies on detecting behavior changes in synthetically modified programs. These mutated versions are highly unlikely to appear in pretraining corpora, providing a more reliable evaluation of CANDOR.
@@
Conclusion Validity. A potential threat is the randomness in LLM outputs, which may affect the consistency of results. To mitigate this, we repeat all experiments three times and apply statistical testing to assess the significance of the results across programs.
External Validity. One threat is that our evaluation is conducted on method-level datasets, and it is unclear how well CANDOR generalizes to more complex, project-level programs that involve dependencies. However, CANDOR is a multi-agent framework, which can be easily extended by introducing dedicated agents to handling dependencies. We leave this exploration for future work. Moreover, the strong performance of CANDOR on the LeetCodeJava dataset, which contains problems of varying difficulty, suggests that it is robust across diverse testing scenarios.
7 Related Work
Traditional methods rely on search-based and randomized techniques to generate test cases for Java programs Fraser and Arcuri (2011b); Rojas et al. (2017); McMinn (2004); Fraser and Arcuri (2011a); Pacheco and Ernst (2007). Random testing tools, such as Randoop Pacheco and Ernst (2007), generate tests by executing sequences of randomly selected method calls. These approaches generate test inputs and regression oracles derived from the current implementation, which are effective for detecting behavioral changes across software versions but limited in validating functional correctness against intended specifications. This limitation arises because regression oracles assume the existing code is correct, potentially reinforcing bugs rather than detecting them. EvoSuite Fraser and Arcuri (2011a), which is the SOTA tool, uses genetic algorithms to systematically generate test suites for Java programs, focusing on maximizing code coverage.
To go beyond regression oracles, a few approaches have been proposed. TOGADinella et al. (2022) introduces a template-based approach for generating functional oracles by leveraging documentation and specification cues. More recently, TOGLLHossain and Dwyer (2024) combines EvoSuite-generated test prefixes with LLM-generated oracles through fine-tuned models. However, TOGLL depends on test prefixes generated by EvoSuite and requires extensive fine-tuning of large models, which can be resource-intensive and less adaptable to new domains. LLM-Empirical Siddiq et al. (2024) offers an alternative by using prompt-engineering to generate JUnit tests based on natural language descriptions, without fine-tuning. It focuses on generating both test prefixes and oracles using only prompt engineering. CANDOR follows this line of research but introduces several key innovations, including the design of multiple specialized agents, the panel discussion to reduce hallucination and the dual-LLM pipeline of to extract structured information from the verbose output of reasoning LLMs. While recent advances in multi-agent LLM systems have achieved success in various software engineering domains (e.g. debugging Lee et al. (2024), fault localization Qin et al. (2024) and program repair Bouzenia et al. (2024)), their application to software testing remains underexplored. CANDOR is, to the best of our knowledge, is the first multi-agent, end-to-end JUnit test generator, laying the groundwork for future research in agent-based automated testing.
8 Conclusion
@@
In this work, we presented CANDOR, a novel multi-agent, end-to-end framework for generating high-quality unit tests with accurate oracles. CANDOR leverages a panel discussion to mitigate hallucinations and employs a dual-LLM pipeline to reduce overthinking during reasoning.
@@
Our extensive evaluation on HumanEvalJava and LeetCode demonstrates that CANDOR achieves better performance than EvoSuite in terms of line coverage and mutation score, while producing accurate specification-based oracles, which EvoSuite cannot. Ablation studies highlight the critical roles of key agents, including the Planner for enhancing test prefix quality and the Requirement Engineer along with the panel discussion for improving oracle correctness. For future work, we plan to explore project-level test generation and investigate more advanced variants of panel discussions to further reduce hallucination.
Acknowledgment
This work was supported by a research grant from Huawei Technologies and Research Ireland grant 13/RC/209. Lionel C. Briand’s contribution was partially funded by the Research Chair and Discovery Grant programs of the Natural Sciences and Engineering Research Council of Canada (NSERC).

### v6

@@ 6.2. Threats to validity
Construct Validity. We evaluate CANDOR using code coverage, mutation score, and oracle correctness. While real bug detection on datasets like Defects4J is another useful metric, we did not include it because such project-level datasets are beyond the scope of this work, as discussed earlier. Instead, we use the mutation score as a proxy for bug-finding capability.
@@
Internal Validity. Our results may be influenced by the choice of LLMs used in the framework. We use LLaMA 3.1 70B as the basic LLM and DeepSeek R1 as the reasoning LLM, selected based on preliminary experiments for their stability and open-source availability. While other popular models, such as GPT, Grok, and Mistral, exist, exhaustively evaluating all LLMs is beyond the scope of this work. The choice of LLMs may affect performance, but we aim to demonstrate the potential effectiveness of our method rather than benchmark specific models, which might indeed yield even better results. Another potential threat is data leakage, where our datasets may be included in the LLMs’ pretraining data, leading to inflated performance. To mitigate this, we evaluate the mutation score, which detects behavioral changes in synthetically modified programs. These mutated versions are highly unlikely to appear in pretraining corpora, providing a more reliable evaluation of CANDOR.
@@
Conclusion Validity. A potential threat is the randomness in LLM outputs, which may affect the consistency of results. Specifically, LLMs can produce different outputs even when given identical inputs. To mitigate this, we repeat all experiments three times, compute the average, and apply statistical testing to assess the significance of differences in averages across programs.
External Validity. One threat is that our evaluation is conducted on only two datasets (i.e., HumanEvalJava and LeetCodeJava). It is unclear how well CANDOR and the baselines (EvoSuite, LLM-Empirical, and TOGLL) generalize to other datasets, especially those that involve complex programs with dependencies on user-defined or external classes. However, HumanEvalJava is a representative, commonly used benchmark dataset in recent studies (jiang2023impact; Siddiq2024UsingStudy; chen2024chatunitest; yuan; kang2025explainable; li2024hybrid; endres2024can; tian2025fixing), while we created LeetCodeJava to demonstrate the effectiveness of CANDOR on programs with higher and varying complexity. Moreover, our experimental results show that automated test generation, especially oracle generation, remains unresolved by existing approaches on these datasets. Our work thus represents an important first step towards generating tests for more complex programs. In fact, CANDOR is a multi-agent framework, which can be readily extended by introducing a dedicated agent to handle dependencies. Such an agent would retrieve definitions of user-defined or external classes, augmenting the test generation process when complex dependencies are involved. We leave this exploration for future work. Last, the strong performance of CANDOR on the LeetCodeJava dataset, which contains problems of varying difficulty, suggests that it is robust across diverse testing scenarios.
7. Related Work
We report prior work on test prefix generation and specification-based test oracle generation in Sections 7.1 and 7.2, respectively.
7.1. Test Prefix Generation
Early works in what is commonly referred to as test case generation are, in fact, more accurately characterized as test prefix generation accompanied by regression oracles. These works in test prefix generation primarily aim to automate manual testing by generating test prefixes with high code coverage. Traditional techniques in this area include fuzzing (Miller1990AnUtilities; Fioraldi2023DissectingEvaluation), feedback-directed random test generation (Csallner2004JCrasher:Java; Pacheco2007Randoop:Java; Pacheco2008FindingTesting; Selakovic2018TestLanguages; Arteca2022Nessie:Callbacks), dynamic symbolic execution (Godefroid2005DART:Testing; Sen2005CUTE:C; Cadar2008EXE:Death; Tillmann2014TransferringDigger), and search/evolutionary algorithm-based approaches (Fraser2011EvolutionarySuites; Fraser2011EvoSuite:Software; Pacheco2007Randoop:Java). Among these approaches, EvoSuite has been extensively studied and applied to more than a hundred open-source software projects and several industrial systems, identifying thousands of potential bugs. However, as noted by zhang2025large, traditional approaches such as EvoSuite often struggle to produce meaningful, human-readable test cases, despite their remarkable code coverage. This is primarily due to their limited understanding of the semantics of the focal methods. To alleviate such issues, recent research has turned to LLMs to produce more practical test prefixes with high readability and coverage.
Tufano2021UnitContext introduced the first LLM-based unit test generation approach AthenaTest, which formulates the task as a sequence-to-sequence learning problem. AthenaTest first denoises the pre-training of LLMs on a large Java corpus and then performs supervised fine-tuning for the downstream task of test generation. This work inspired subsequent research on fine-tuning LLMs for test generation (ALAGARSAMY2024107565; catlm; unitysyn). Additionally, researchers have explored various strategies to improve LLM training for test case generation, including reinforcement learning (reinforce), domain adaptation (domain), and data augmentation (aug).
Alternatively, a myriad of works perform prompt engineering using off-the-shelf LLMs, following a generation-and-refinement paradigm, where initial test cases are first generated based on prompts and then iteratively refined using dynamic execution feedback (e.g., code coverage report and failure information) (chen2024chatunitest; gu2024testart; ni2024casmodatest; altmayer2025coverup; codeaware; Schafer2024AnGeneration; wang2024hits; yuan; zhang2025large). To further improve the quality of test prefixes, researchers propose a wide range of strategies to incorporate LLMs with valuable contextual information, including mutation testing (dakhel2024effective), method slicing (wang2024hits), demonstration retrieval (zhang2024llm), defect detection (yin2025you), and program analysis (yang2024enhancing).
Despite the rapid growth in LLM-based test generation research, tang2024chatgpt reported that EvoSuite remains the SOTA approach in generating test prefixes with high code coverage. CANDOR builds on prompt engineering-based test generation and makes the following contributions: (1) CANDOR is the first multi-agent LLM-based test generator. CANDOR lays the foundation for multi-agent frameworks in test generation by proposing a generic, effective framework with carefully designed agents, including Planner, Tester, and Inspector. (2) CANDOR presents a breakthrough in LLM-based test prefix generation, reaching comparable code coverage and significantly better mutation rates than evo.
7.2. Test Oracle Generation
Test oracle generation requires an accurate understanding of program semantics and requirements, posing significant challenges for early automated testing tools (chen2024chatunitest; gu2024testart; ni2024casmodatest; altmayer2025coverup; codeaware; Fraser2011EvoSuite:Software; Pacheco2007Randoop:Java). These tools circumvent these challenges by generating regression oracles, while specification-based oracle generation has been underexplored. Regression oracles are derived from the current implementation and are effective for detecting behavioral changes across software versions; however, they are limited in validating functional correctness against intended specifications. This limitation arises because regression oracles assume the existing code is correct, potentially masking bugs rather than detecting them.
However, going beyond regression oracles and generating specification-based oracles is challenging because it requires a deep understanding of functional requirements, edge cases, and expected outcomes (binta2024doc2oracle). With the advent of LLMs, researchers have begun to explore their potential for generating accurate specification-based oracles, using both prompt engineering and fine-tuning strategies. Dinella2022Toga:Generation introduces TOGA, a transformer-based approach for inferring specification-based oracles from the context of the focal method. TOGA is the first LLM-based test oracle generator powered by a fine-tuned CodeBERT LLM. Inspired by this work, several studies have proposed generating specification-based oracles using either FT or PE strategies (chatassert; Hossain2024TOGLL:LLMs; chatassert; catlm; endres2024can; Siddiq2024UsingStudy). Siddiq2024UsingStudy proposed LLM-Empirical, a solution using prompt engineering to generate JUnit tests based on natural language descriptions, without fine-tuning. It focuses on generating both test prefixes and oracles using only prompt engineering. zhang2025exploring conduct the first comprehensive study on fine-tuning various LLMs for oracle generation across two benchmarks, five LLMs, and two metrics. Building on this study, zhang2025improving introduced RetriGen, a retrieval-augmented oracle generation approach that incorporates a novel hybrid assertion retriever into the oracle generation process. RetriGen retrieves most relevant assertions from external codebases by leveraging both lexical and semantic similarity. TOGLL(Hossain2024TOGLL:LLMs) combines EvoSuite-generated test prefixes with LLM-generated oracles through fine-tuned models, achieving SOTA performance in oracle generation for Java programs. However, TOGLL depends on test prefixes generated by EvoSuite and requires the extensive fine-tuning of large models, which can be resource-intensive and less adaptable to new domains.
CANDOR also relies on prompt engineering-based oracle generation but introduces several key innovations, including the design of multiple specialized agents, the panel discussion to reduce hallucination, and the dual-LLM pipeline to extract structured information from the verbose output of reasoning LLMs.
8. Conclusion
@@
In this work, we presented CANDOR, a novel multi-agent, end-to-end framework for generating high-quality unit tests with accurate oracles. CANDOR does not require fine-tuning or the use of external tools but leverages a panel discussion to mitigate hallucinations and employs a dual-LLM pipeline to reduce overthinking during reasoning.
@@
Extensive evaluation on two benchmarks, HumanEvalJava and LeetCode, demonstrates that CANDOR achieves performance comparable to EvoSuite in terms of line coverage and branch coverage, while substantially outperforming EvoSuite in mutation score. Additionally, CANDOR produces accurate specification-based oracles that outperform the state-of-the-art oracle generator TOGLL by a large margin ( percentage points). Ablation studies highlight the critical roles of key agents: the Planner for enhancing test prefix quality, and the Requirement Engineer together with the panel discussion for improving oracle correctness. For future work, we plan to extend CANDOR to project-level test generation and explore more advanced panel discussion strategies to further reduce hallucinations.
Appendix
CANDOR with Alternative LLMs
| Dataset | Metric | Llama | CodeLlama | Mistral |
| HumanEvalJava | Line | 0.991 | 0.991 | 0.840* |
| Branch | 0.970 | 0.967 | 0.839* | |
| Mutation | 0.980 (2384/2443) | 0.980 (2390/2443) | 0.796* (1945/2443) | |
| Oracle | 0.910 | 0.912 | 0.861* | |
| Leetcode-Medium | Line | 0.990 | 0.990 | 0.914* |
| Branch | 0.949 | 0.950 | 0.943* | |
| Mutation | 0.939 (574/611) | 0.948 (579/611) | 0.839* (513/611) | |
| Oracle | 0.930 | 0.919 | 0.904* | |
| Leetcode-Hard | Line | 0.989 | 0.990 | 0.876* |
| Branch | 0.980 | 0.981 | 0.871* | |
| Mutation | 0.937 (805/859) | 0.930 (799/859) | 0.905* (777/859) | |
| Oracle | 0.894 | 0.869* | 0.848* |
Table 3 reports the experimental results of CANDOR using alternative LLMs. We observe that Mistral 22B has the lowest effectiveness among the three models, resulting in substantially lower line coverage, branch coverage, mutation scores, and oracle correctness. Wilcoxon Signed Rank tests confirm that all differences between Mistral 22B and Llama 70B are statistically significant, with p-values below . This outcome is expected, as Mistral 22B has considerably fewer parameters than Llama 70B and CodeLlama 70B, making it less capable across tasks, including test generation.
In contrast, CodeLlama 70B and Llama 70B yield comparable results. For test prefix generation, the largest difference between the two models is a 0.09 gap in mutation score on the LeetCode-Medium dataset. However, Wilcoxon Signed Rank tests show that none of these differences are statistically significant. For oracle generation, Llama 70B achieves slightly higher correctness on LeetCode-Medium () and LeetCode-Hard (), while performing slightly worse on HumanEvalJava (). The differences between HumanEvalJava and LeetCode-Medium are not statistically significant, whereas the difference between LeetCode-Hard and HumanEvalJava is. Overall, Llama 70B performs marginally better than CodeLlama 70B. For this reason, we select Llama 70B as the default model for CANDOR and report the corresponding results in the main manuscript.